If the distance between earth and sun were reduced to half of its present value , then the number of days in one year would have been


a) 65


b) 129


c) 183


d) 730

Grade:11

4 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Arnab,

 

The ratio of the time periods would be {new is say T1, actual is say T2}

 

Note that T is directly proportional to √(R3)

 

T1/T2 = √([R/2]3)/√(R3) = 1/(2√2)

 

Hence T1 = T2/2√2 = 365/2√2

 

Which is approx 129 days.

Option (B)

 

Hope that helps.

 

Best Regards,

Ashwin (IIT Madras).

nikhil manali
12 years ago

no its wrong option as it is approx to 183 days

nikhil manali
12 years ago

yup it is wright ans

 

Ritik
6 years ago
Yah.. ans is B . It can be solved by kepler`s third law . Which is T^2 proportional to R^3. Where To is period of revolution and R is distance.